Is each ordered pair a soulution of 4x-3y=6
(-2,4) (1,-2) (4,-8)
cant figure out how to solve problem i only have examples for problems with one ex: 4x-y=6

Answer by quentin252(19)

simply plug in first digit for x
and second digit to y
so... (-2,4) -->> 4(-2)-3(4) = -8 - 12 = -20 not equal to 6
(1,-2) -->> 4(1)-3(-2) = 4 -(-6) = 4+6 = 10
and so on...(4,-8)-->>40